OpenStudy (some.random.cool.kid):

can someone give me examples of addition subtraction and multiply divide polynomials and solve them with me?

OpenStudy (some.random.cool.kid):

@dtan5457

OpenStudy (some.random.cool.kid):

@SolomonZelman

OpenStudy (some.random.cool.kid):

@ganeshie8

OpenStudy (some.random.cool.kid):

@TheSmartOne

OpenStudy (some.random.cool.kid):

@undeadknight26

OpenStudy (some.random.cool.kid):

@Compassionate

OpenStudy (some.random.cool.kid):

@sleepyhead314

OpenStudy (sleepyhead314):

(x^2+x+2)+(2x^2+3x+7) = ? you mean like that?

OpenStudy (some.random.cool.kid):

yes

OpenStudy (some.random.cool.kid):

except not as many steps.

OpenStudy (sleepyhead314):

(5x-7) + (2x + 3) = ?

OpenStudy (some.random.cool.kid):

more like this : 8x² – 5x – 2 + 7x – x² + 3 I solved this one already but this is just an example

OpenStudy (some.random.cool.kid):

oh yeah thats a great one sleepy

OpenStudy (sleepyhead314):

tbh khanacademy probably has a problem generator instead of me trying to come up with random questions out of my head xD

OpenStudy (some.random.cool.kid):

that one works though thats actually perfect.

OpenStudy (some.random.cool.kid):

how would i solve this (5x-7) + (2x + 3) except with it like (5x-7+ 7x^3) + (2x + 3 + 5x^2)

OpenStudy (sleepyhead314):

just add up thing things that have the same letter things 5x + 2x = 7x

OpenStudy (some.random.cool.kid):

for instance i have this one to do ... (2𝑥^2 +𝑥+5 +(3𝑥^2 +4𝑥+1)

OpenStudy (some.random.cool.kid):

yeah that one is easy this one is a it tougher.

OpenStudy (sleepyhead314):

"add like terms" 2x^2 + 3x^2 = 5x^2

OpenStudy (some.random.cool.kid):

and 5 + 1 is 6 so?

OpenStudy (some.random.cool.kid):

im guesssing 5x^2 + 6 + 4x= ?

OpenStudy (some.random.cool.kid):

alright final one

OpenStudy (sleepyhead314):

not quite

OpenStudy (sleepyhead314):

don't forget the 4x + x = 5x

OpenStudy (some.random.cool.kid):

ok

OpenStudy (some.random.cool.kid):

i got it now but can we do one with some minuses in it

OpenStudy (sleepyhead314):

sure :) do you have one or would you like for me to make one up?

OpenStudy (some.random.cool.kid):

can you make one please?

OpenStudy (sleepyhead314):

4x^2 - 7x + 1 - x^2 + 3x - 9

OpenStudy (some.random.cool.kid):

7x + 3x = 10x

OpenStudy (sleepyhead314):

see the negative?

OpenStudy (sleepyhead314):

-7x + 3x

OpenStudy (some.random.cool.kid):

oh so 4x

OpenStudy (some.random.cool.kid):

or -4x

OpenStudy (sleepyhead314):

negative still -4x

OpenStudy (some.random.cool.kid):

i need to do multiply now here is one

OpenStudy (some.random.cool.kid):

4x^2 + 2x -1) x ( 3x^3 – x + 4)

OpenStudy (some.random.cool.kid):

you there @sleepyhead314 ?

OpenStudy (sleepyhead314):

(a + b - c)*(x - y + z) = a*x + a*(-y) + a*z + b*x + b*(-y) + b*z + (-c)*x + (-c)*(-y) + (-c)*z

OpenStudy (some.random.cool.kid):

how does this help?

OpenStudy (sleepyhead314):

sorry, that came out messy >,<it's just a longer distributive property